OpenStudy (anonymous):

9x^2 + 16x = 0 how to complete the square

OpenStudy (anonymous):

x(9x+16)=0 9x=-16 x=-16/9 and x=0

OpenStudy (solomonzelman):

what number added to both sides would make the left side a perfect square?

OpenStudy (anonymous):

16?

OpenStudy (solomonzelman):

I never thought of it formulaicly, but \[(3x+\text{_})^2\]you need to fill in the blank so that the middle term would be =16

OpenStudy (anonymous):

4^2 = 16

OpenStudy (anonymous):

can you show me step by step

OpenStudy (solomonzelman):

well also times 3 and times 2

OpenStudy (solomonzelman):

\[3 \times x \times 2~=16\]

OpenStudy (solomonzelman):

as \[(a+b)^2=a^2+\color{red} {2ab}+b^2\]\[2(3x )~~~~is~~~~your~~~~2ab,~~~~o r~~~~th e~~~~m iddle~~~~term\]

OpenStudy (solomonzelman):

Do you see what i am doing?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\[(3x^2~+~\text{__}~)~\]fill in blank.

OpenStudy (solomonzelman):

I'll tell you, alright....

OpenStudy (solomonzelman):

\[\frac{8}{3}\]

OpenStudy (solomonzelman):

\[(3x+\frac{8}{3})^2=0+(\frac{8}{3})^2\]

OpenStudy (solomonzelman):

can you do it from here?